The Northern Co. manufactures office chairs. The following information is available for June: Selling price per chair $150 Variable cost per chair 90 Total fixed costs 99,000 If the company sells 1,500 chairs in the month of June, the margin of safety in units is _____. (Round-off the answer to the nearest whole number.)

a.-179 chairs b.-150 chairs c.-159 chairs d.-169 chairs
Answer #1

Contribution margin=Sales-Variable cost  

=150-90

=$60 per chair

Breakeven=Fixed expenses/Contribution margin

=99000/60

=1650 chairs

Margin of safety=Total sales-Breakeven sales

=1500-1650

=-150 chairs